Consider the probability distribution ?(?) = ??n, 0 ≤ ? ≤ 1 for a positive integer...

  1. Consider the probability distribution ?(?) = ??n, 0 ≤ ? ≤ 1 for a positive integer ?.

    1. Derive an expression for the constant ?, to normalize ?(?).

    2. Compute the average 〈?〉 as a function of ?.

    3. Compute the expectation value of the second moment.

    4. Compute the variance as a function of ?.
Answer #1

Given Pdf is- p(2)= aren, osasi L (1) To obtain (9) I praydn = 1 - Jaandaal a = 1 → a= nti thus , p(a)= (n+1) n o ToLas (2) <(3) <^2) = E (a) = (*_*(ntvmdn ht) nt3 and order moment (4) variance = Ln²> - [<ny] ht3 nta, = (n+1)* {(n+2)2-(n+1)(n+3)? (n+